OpenStudy (anonymous):

sqrt(9x+4) integrate from 0 to 1

hartnn (hartnn):

try to make a u substitution, u = 9x+4 du =... ?

OpenStudy (anonymous):

du=9??

hartnn (hartnn):

thats correct! so whats your new integral ?

hartnn (hartnn):

don't forget to change the limits too when x = 0, u =....? when x = 1, u =... ?

OpenStudy (anonymous):

when x=0 u=4 when x=1 u=13

OpenStudy (anonymous):

\[\sqrt{u}\]

OpenStudy (anonymous):

or (u)^(1/2)

hartnn (hartnn):

correct! so you can proceed further ? integral of sqrt u is ?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

2u?

hartnn (hartnn):

no, \(\int x^n dx = x^{n+1}/(n+1)+c\) so, for integral of u^1/2, plug in n = 1/2 there

OpenStudy (anonymous):

\[(2/3)u^(2/3)\]

OpenStudy (anonymous):

+c

hartnn (hartnn):

1/2 +1 is actually 3/2 so, \(\Large \dfrac{u^{3/2}}{(3/2)} = \dfrac{2 u^{3/2}}{3}+c\)

hartnn (hartnn):

and then apply your limits for u

OpenStudy (anonymous):

duh sorry about that

OpenStudy (anonymous):

Thanks so much

OpenStudy (whpalmer4):

wait, haven't you got the wrong fraction multiplying \(u^{3/2}\)?

OpenStudy (whpalmer4):

you're missing the \(9\) from \(du = 9\, dx\) I believe...

OpenStudy (whpalmer4):

\[du = 9\,dx\]\[dx = \frac{1}{9}du\]

hartnn (hartnn):

oh yes! thanks for pointing that out @whpalmer4 :)

OpenStudy (whpalmer4):

then you have (after moving the constant fraction outside the integral) \[\frac{1}{9}\int u^{\frac{1}{2}}\,du\] (too lazy to put on the limits)

OpenStudy (anonymous):

\[\frac{ 1}{ 9} (\frac{ 2(13)^{\frac{ 2 }{ 3 }}}{ 3 })-(\frac{ 2(4)^{\frac{ 2 }{ 3 }}}{ 3 })\]

OpenStudy (anonymous):

is this correct??

OpenStudy (anonymous):

Then evaluate?

OpenStudy (whpalmer4):

almost, you've got the exponent fractions upside down. I'd simplify (after fixing exponents) to \[\frac{2}{27}(13^{\frac{3}{2}}-4^{\frac{3}{2}})\]rather than do each of those terms as fractions

OpenStudy (whpalmer4):

(still evaluate it, of course!)

OpenStudy (whpalmer4):

I do integrals just often enough to remember where I usually make mistakes doing integrals :-)

OpenStudy (anonymous):

Thank you for your help on this I will be back with a answer.

OpenStudy (anonymous):

2/27(13sqrt(13)-8)